OpenStudy (tmostiller):

find the x- and y- intercept of the line. -7/5x-4y=7

OpenStudy (08surya):

put x=0 and calculate y put y=0 and calculate x

OpenStudy (aveline):

Plug in 0 for x then 0 for y. -7/5(0)-4y=7 -4y=7 (0, -7/4) -7/5x-4(0)=7 -7/5x=7 x=-5 (-5, 0)

OpenStudy (tmostiller):

I have to choose from a. x-intercept is -7/4; y-intercept is 5 b. x-intercept is 5; y-intercept is 7/4 c. x-intercept is 5; y-intercept is -7/4 d. x-intercept is -5; y-intercept is -7/4 this is the only one I just don't grasp. Thanks

OpenStudy (tmostiller):

duh, it's D right?

OpenStudy (am16):

To calculate y intercept put 0 in place of x in the equation: then it becomes=> -4y=7 so y= -7/4 so point is (0,-7/4) To calculate x intercept put 0 in place of y in the equation: then it becomes=> (-7/5)x=7 so x= -(7*5)/7 => x= -5 so point is (-5,0)